1. How do I find the equation of the line passing through the point (1,1) and parallel to y=5x-6?
2. How do I find the equation of the line passing through the point (1,5) and parallel to 4x+7y=18?
Hello, I am having a very hard time with this coordinate geometry problem. Can anyone please help? Thanks so much!

Answer by sudhanshu_kmr(1152)


1. slope of line = 5
equation of line,
(y-1) = 5 (x-1)
=> y -1 = 5x -5
=> 5x -y = 4
you are welcome to contact me for any help........